Fairfield Learning is a retailer focused on education supplies. The company has a book value of $28.88 per share. Fairfield Learning has a PB ratio of 6.23 and the education supplies industry PB ratio is 5.31. Assuming that comparable industry companies are priced correctly the intrinsic value of Fairfield Learning's equity per share is:

Undervalued $26.57 per share

Overvalued $26.57 per share

Priced correctly

Overvalued by $29.17 per share
Answer #1

Price to Book Ratio (PB) is considered as an important ratio to find whether the share of company is Overvalued or Undervalued. In the above question

Book Value = $28.88

PB ratio of company = 6.23

So, Price = 6.23*28.88 = $179.92

The Industry PB ratio = 5.31

So, Price should be = 5.31*28.88 = $153.35

As the PB ratio of Industry < PB ratio of Company , therefor Share is Overvalued by 179.92-153.35 = $26.57